OpenStudy (djbreezy):

Which is an equation of given line in standard form? A. -11x+6y=-13 B. -6x+11y=13 C. -6x+11y =-13 D. -6x+7y=-17

OpenStudy (djbreezy):

Line has (-4,-1) and (1 1/2,2)

sam (.sam.):

First, find the slope by using the formula \[m=\frac{y-y_1}{x-x_1}\] Once you've got the slope, simply pick one of the given coordinates, say (-4,-1), and plug it into \[(y-y_1)=m(x-x_1)\] Where \(y_1\) is -1, and \(x_1\) is -4.

OpenStudy (djbreezy):

is it B

OpenStudy (karim728):

did you find the slop?

OpenStudy (djbreezy):

@ToxicWaste

OpenStudy (dallasrichards16):

It is B

OpenStudy (toxicwaste):

sorry @DJBreezy but im not really good at math

OpenStudy (djbreezy):

kk

OpenStudy (neonumbrella5115):

Do you still need help? @DJBreezy

OpenStudy (djbreezy):

yaahh

OpenStudy (neonumbrella5115):

Okay! To find the slope we can use the equation about and subtract\[\frac{ 2--2 }{ 1.5--4 }\]

OpenStudy (neonumbrella5115):

And when you solve that expression, what do you get? @DJBreezy

OpenStudy (djbreezy):

U get 0.72727272

OpenStudy (mathstudent55):

@neonumbrella5115 Aren't the y-coordinates of the two points 2 and -1? Check your numerator above.

OpenStudy (neonumbrella5115):

My mistake! The real equation is \[\frac{ 2--1 }{1.5--4 }\] Thanks!! @mathstudent55

OpenStudy (mathstudent55):

No. Your numerator needed fixing, and you did. Your denominator was correct. Now it's missing the - of -4.

OpenStudy (mathstudent55):

Oh, you fixed it again. Now you're good.

OpenStudy (neonumbrella5115):

@DJBreezy What do you get when you solve the now-correct equation? xp

OpenStudy (djbreezy):

0.54545454

OpenStudy (mathstudent55):

Your number for the slope is correct, but leave it as a fraction.

OpenStudy (neonumbrella5115):

Now we'll write the equation in slope intercept form. so y=mx+b However, we don't know the b value yet so we will have to solve for that. First plug the slope we just found in for m, and then you will plug in the x and y values from one of the coordinates in the original equation.

OpenStudy (djbreezy):

3/1.9

OpenStudy (mathstudent55):

\(m = \dfrac{ 2-(-1) }{1.5-(-4) } = \dfrac{3}{5.5} = \dfrac{6}{11}\) The slope is 6/11. Now use the equation of a line given by @.Sam. above. Plug in the slope we found above, 6/11, and the coordinates -1 and -4. That will give you the equation of the line. Then you need to put it in standard form: \(Ax + By = C\)

OpenStudy (mathstudent55):